give me the answer and its solution:
m=1/2 (one half)
PTa=(-2,-4)
PTb=(k,4)
Find:K=?
[hint: Y2-Y1/X2-X1 (Y2 minus Y1 over X2 minus X1)
Then graph it please.
Thats what my teacher told me..

Answer by stanbon(75887)

m=1/2 (one half)
PTa=(-2,-4)
PTb=(k,4)
Find:K=?
------
slope = (-4-4)/(-2-k) = 1/2
---
-8/(-2-k) = 1/2
Cross-multiply:
-16 = -2-k
k = 14
===========
Cheers,
Stan H.
===========